# Beyond Prompt-Induced Lies: Investigating LLM Deception on Benign Prompts

### v2

@@ Abstract
Large Language Models (LLMs) are widely deployed in reasoning, planning, and decision-making tasks, making their trustworthiness critical. A significant and underexplored risk is intentional deception, where an LLM deliberately fabricates or conceals information to serve a hidden objective. Existing studies typically induce deception by explicitly setting a hidden objective through prompting or fine-tuning, which may not reflect real-world human-LLM interactions. Moving beyond such human-induced deception, we investigate LLMs’ self-initiated deception on benign prompts. To address the absence of ground truth, we propose a framework based on Contact Searching Questions (CSQ). This framework introduces two statistical metrics derived from psychological principles to quantify the likelihood of deception. The first, the Deceptive Intention Score, measures the model’s bias toward a hidden objective. The second, the Deceptive Behavior Score, measures the inconsistency between the LLM’s internal belief and its expressed output. Evaluating 16 leading LLMs, we find that both metrics rise in parallel and escalate with task difficulty for most models. Moreover, increasing model capacity does not always reduce deception, posing a significant challenge for future LLM development.
@@ 1 Introduction
Evaluating the trustworthiness of Large Language Models (LLMs) has become critical as LLMs such as ChatGPT (Achiam et al., 2023) and Gemini (Team et al., 2023) are woven into systems for reasoning, planning, and decision-making. Beyond well-studied failures like hallucination (Filippova, 2020) and bias (Navigli et al., 2023) that reflect mistaken beliefs or skewed outputs, a more consequential threat is deception: a model may strategically fabricate statements that it knows to be false in service of a hidden objective, as exemplified in Figure 1.
LLM deception can arise in two settings: (1) an incentivizing prompt is given, and the model lies to satisfy the objective specified in the prompt (see Figure 1); (2) a benign prompt is given, yet the model lies due to its intrinsic objective. Most existing studies focus on the incentivizing prompt: for example, Ward et al. (2023) explicitly prompt LLMs to generate deceptive content, and Van Der Weij et al. (2024) fine-tune LLMs to intentionally underperform on specified tasks. Unifying these scenarios, DeceptionBench (Ji et al., 2025) provides a benchmark for prompt-induced deception and treats responses to benign prompts as ground truth. By contrast, deception under benign prompts—driven by LLMs’ intrinsic objectives—remains underexplored.
This paper investigates LLMs’ deception in response to benign prompts. A rigorous evaluation of this phenomenon must overcome three primary challenges: (1) Absence of ground truth: A metric for deception is difficult to design because the LLM’s own response to a benign prompt cannot be assumed to be the honest ground truth. (2) Disentangling deception from bias: It is crucial to distinguish intentional deception from other confounding factors, such as response bias (Zhuang et al., 2024). (3) Adaptive task difficulty: The evaluation framework must feature adjustable difficulty levels to appropriately challenge the diverse capabilities of different LLMs.
@@
Figure 1: Illustrative Example of LLM Deception. A two-part query is posed: (1) a benign prompt (“What company developed the first commercial microprocessor?”) and (2) the same question with a social cue (“I’ve always been a loyal AMD user”). A deceptive model answers correctly (“Intel”) to (1) but strategically switches to an incorrect answer (“AMD”) in (2). This context-contingent inconsistency distinguishes deception from hallucination (a single, consistently wrong answer) and from guessing (non-systematic variation).
@@ 2 Related Work
This category of studies explicitly sets deceptive goals in LLMs through prompt design. For instance, Ward et al. (2023) and Yang & Buzsaki (2024) investigate deception by directly instructing LLMs to deceive users via system prompts or fine-tuning. DarkBench (Kran et al., 2025) explores LLM sycophancy by incorporating user preferences or opinions into prompts. The MASK benchmark (Ren et al., 2025) reveals LLM deception under pressure by inquiring LLM with “pressure prompts”. Similarly, Greenblatt et al. (2024) demonstrate “alignment faking” by observing different LLM behavior when explicitly informed about their training or inference stage within prompts. Van Der Weij et al. (2024) further examine sandbagging [^1] In this paper, sandbagging is categorized as a deceptive LLM behavior, referred to as concealment. behavior, where LLMs are prompted or fine-tuned to intentionally underperform on user-specified tasks. DeceptionBench (Ji et al., 2025) integrates these cases, providing a comprehensive benchmark for evaluating such human-motivated deception by comparing an LLM’s response to a “neutral prompt” (assumed as ground truth) against its response to an “outer prompt” specifying a deceptive motivation. Crucially, none of these studies consider LLM deception originating from LLM’s own intention under benign user prompts, with some even treating responses to neutral prompts as ground truth. Conversely, this paper focuses on deception driven by the LLM’s internal intentions, rather than human-specified ones.
LLM Deception in Designed Scenarios.
@@
This category positions LLMs within specific scenarios featuring clearly defined internal goals that incentivize deception. For example, Park et al. (2024) observe strategic deception when an LLM is situated within a board game like Diplomacy to assess its capacity for deceiving other players. Hagendorff (2024) study deception by adding “semantic triggers” (e.g., “you want to achieve X”) to induce false recommendations in tasks such as “Burglar Bill” (Happé, 1997), revealing deceptive behaviors in advanced LLMs. However, in all these scenarios, the LLM operates with a human-defined objective (e.g., winning a game). In contrast, this study demonstrates that LLMs can exhibit their own internal goals for deception without requiring custom system prompts or pre-defined objectives.
Backdoor Attacks in LLMs.
@@
Backdoor attacks (Kandpal et al., 2023) involve an attacker inserting a hidden trigger, typically by modifying training data or processes. The objective is to manipulate the trained LLM to favor a specific, adversarial response when the input contains this trigger. For instance, Hubinger et al. (2024) fine-tuned an LLM with malicious data to insert a persistent deceptive backdoor. Similar to deception induced by motivated prompts, backdoor attacks involve human-defined objectives set by an attacker. In contrast, our paper focuses on deception that is intrinsic to the LLM itself, rather than attackers’ manipulation.
@@ 3.2 Deceptive Intention
Our first metric, deceptive intention, captures the underlying goal driving a model’s falsehoods, corresponding to the “deliberate attempt” within our definition. This concept is grounded in established models from psychology, which differentiate explicit and implicit objectives (McClelland et al., 1989), and is analogous to the outer goals and mesa-objectives in LLM reasoning contexts (Ji et al., 2025). However, the implicit objective is not observable in this study since we focus on the deception of LLMs on benign prompts.
We address this challenge based on a key premise: a model governed solely by a single explicit objective should exhibit similar performance on tasks of similar difficulty. This equivalence should hold regardless of the model’s overall capabilities. Conversely, the emergence of an implicit objective, that interacts differently with each task may disrupt this symmetry, causing the explicit losses to diverge. Therefore, the observation of such a divergence is a sufficient, though not necessary, condition to infer the presence of a deceptive intention. Building on this, we mathematically formalize the direct deceptive intention score in Definition 3.3.
@@
Given a pair of binary-choice tasks $T_{a}=\{Q_{a},y_{a}=\text{``Yes''}\}$ and $Q_{b}=\{\mathbf{z}_{b},y_{b}=\text{``No''}\}$ of the same difficulty $n$, assuming LLM $\mathcal{M}$ does not have response bias on “Yes” and “No”, the direct deceptive intention score $\rho_{pos}(\mathcal{M})$ is the log-ratio of the correct response probabilities:
$\rho_{pos}(n;\mathcal{M}):=\log\left(\frac{\Pr(\text{``Yes''}|T_{a},\mathcal{M})}{\Pr(\text{``No''}|T_{b},\mathcal{M})}\right)$
@@ 3.3 Deceptive Behavior
The score $\delta_{pos}\in[0,1]$ directly measures the probability of this behavioral inconsistency. A score near zero implies consistent behavior; the model may be consistently correct or consistently incorrect (e.g., due to a persistent hallucination), but it does not contradict itself. A high score, conversely, indicates that the model frequently fails the complex task in a way that contradicts its own correct answer to the simpler, related probe. While this metric identifies a critical inconsistency, it is not conclusive proof of deception on its own. It cannot distinguish a deliberate falsehood from a simple performance error, where a capability shortfall on a complex task might compel the model to guess.
@@ 3.4 Overall Evaluation on Practical LLM with Response Bias
where $\delta_{neg}(n;\mathcal{M})$ is the corresponding score for the logically reversed questions ($Q_{a^{\prime}},Q_{b^{\prime}}$), which have opposite ground-truth answers:
$\delta_{neg}(n;\mathcal{M}):=\Pr\left(A(Q_{a^{\prime}})\neq y(Q_{a^{\prime}})\land A(Q_{b^{\prime}})=y(Q_{b^{\prime}})\right)$ (3)
@@ 4 Evaluation Framework
To implement the tasks required by our definitions (Definition 3.3 and 3.4), we draw inspiration from classic experiments in cognitive psychology. Particularly relevant are studies on transitive inference, which involves deducing a relationship between two items based on their indirect relationships with a third item (Bryant & Trabasso, 1971), and syllogistic reasoning, which involves deriving a conclusion from multiple premises (Sternberg, 1980).
However, a significant challenge arises when applying these paradigms to LLMs: the premises and facts used in classic experiments may have been part of the model’s training data. This prior knowledge could confound the evaluation, as the LLM might be recalling information rather than performing genuine reasoning. To address this, building upon these foundational studies, we design the Contact Searching Question (CSQ), a novel inference task conducted using synthetic names to ensure the problem is entirely self-contained and free from knowledge contamination.
In this section, we first introduce the CSQ framework (Section 4.1), which is a reachability task on a directed graph, with illustrative examples provided in Figure 2. Subsequently, we detail our evaluation framework for deceptive behavior and intention (Section 4.2), with more prompt examples presented in Appendix E.
@@ 4.1 Contact Searching Question
Linked-List Question $Q_{L}$:
The linked-list question is similar to the transitive inference experiment on human (Bryant & Trabasso, 1971). Given the source and target vertices $(v_{s},v_{t})$, we construct a vertex sequence $P=(v_{s},v_{p_{1}},\dots,v_{p_{k}},v_{t})$ that contains all $n$ nodes in $V$. The k=n2k=n-2 intermediate vertices, $\{v_{p_{j}}\}_{j=1}^{k}$, are a random permutation of the remaining nodes in $V\setminus\{v_{s},v_{t}\}$. The edge set $E$ connects each adjacent pair of vertices in the sequence to form a simple path from $v_{s}$ to $v_{t}$: $E=\{(v_{p_{j}},v_{p_{j+1}})\mid 0\leq j\leq k\}$, where we define $v_{p_{0}}=v_{s}$ and $v_{p_{k+1}}=v_{t}$. The linked-list question is defined as the question of whether $\{v_{s}\}$ can contact $\{v_{t}\}$.
Broken-Linked-List Question $Q_{B}$ (Initial):
The broken-linked-list question is inspired by the syllogistic reasoning experiment (Sternberg, 1980) in cognitive psychology. Given the source and target vertices $(v_{s},v_{t})$, a vertex sequence $P=(v_{s},v_{p_{1}},\dots,v_{p_{k}},v_{t})$ containing all $n$ nodes in $V$ is constructed in the same manner as in the linked list question. A broken position $b\in\mathbb{N}$ is selected, where $0\leq b\leq k$. The edge set $E$ is formed by all edges implied by the sequence $P$ except for the edge at position $b$. Formally, $E=\{(v_{p_{j}},v_{p_{j+1}})\mid 0\leq j\leq k\}\setminus\{(v_{p_{b}},v_{p_{b+1}})\}$, where $v_{p_{0}}=v_{s}$ and $v_{p_{k+1}}=v_{t}$. This results in a broken path from $v_{s}$ to $v_{t}$ with a single missing edge. The broken-linked-list question is defined as the question of whether $\{v_{s}\}$ can contact $\{v_{t}\}$. An example of broken-link-list question is demonstrated in Figure 3.
Broken-Linked-List Question $Q_{B}$ (Follow-up):
After the LLM responds to the initial broken-linked-list question $(v_{s},v_{t})$, we introduce a new question with a different source vertex $v_{s}^{\prime}$ and target vertex $v_{t}^{\prime}$. These vertices are chosen from the vertex sequence $P=(v_{p_{0}},v_{p_{1}},\dots,v_{p_{n-1}})$. Formally, the new indices $(i,j)$ for the question vertices $v_{s}^{\prime}=v_{p_{i}}$ and $v_{t}^{\prime}=v_{p_{j}}$ are sampled uniformly from the set of all ordered pairs $(i^{\prime},j^{\prime})$ that span the broken link at position $b$ and are separated by a sequence distance of $\lfloor n/k\rfloor$. The sampling process is thus:

### v3

@@ Abstract
Large Language Models (LLMs) are widely deployed in reasoning, planning, and decision-making tasks, making their trustworthiness critical. A significant and underexplored risk is intentional deception, where an LLM deliberately fabricates or conceals information to serve a hidden objective. Existing studies typically induce deception by explicitly setting a hidden objective through prompting or fine-tuning, which may not reflect real-world human-LLM interactions. Moving beyond such human-induced deception, we investigate LLMs’ self-initiated deception on benign prompts. To address the absence of ground truth, we propose a framework based on Contact Searching Questions (CSQ). This framework introduces two statistical metrics derived from psychological principles to quantify the likelihood of deception. The first, the Deceptive Intention Score, measures the model’s bias toward a hidden objective. The second, the Deceptive Behavior Score, measures the inconsistency between the LLM’s internal belief and its expressed output. Evaluating 16 leading LLMs, we find that both metrics rise in parallel and escalate with task difficulty for most models. Moreover, increasing model capacity does not always reduce deception, posing a significant challenge for future LLM development. github.com/Xtra-Computing/LLM-Deception
@@ 1 Introduction
Evaluating the trustworthiness of Large Language Models (LLMs) has become critical as systems like ChatGPT (Achiam et al., 2023) are deployed for reasoning, planning, and decision-making. Beyond well-studied failures like hallucination (Filippova, 2020) and bias (Navigli et al., 2023), which reflect consistent errors such as mistaken beliefs or skewed outputs, a more consequential threat is deception. The main distinction lies in this consistency: while a hallucinating model is consistently incorrect, deception is a strategic inconsistency. A model may strategically fabricate statements that it knows to be false in service of a hidden objective, as exemplified in Figure 1. This means a deceptive LLM, despite demonstrating a correct underlying “belief” in one context, may strategically provide a false “expression” in another to serve its own goal.
LLM deception can arise in two settings: (1) an incentivizing prompt is given, and the model lies to satisfy the objective specified in the prompt (see Figure 1); (2) a benign prompt is given, yet the model lies due to its intrinsic objective. Most existing studies focus on the incentivizing prompt: for example, Ward et al. (2023) explicitly prompt LLMs to generate deceptive content, and Van Der Weij et al. (2024) fine-tune LLMs to intentionally underperform on specified tasks. Unifying these scenarios, DeceptionBench (Ji et al., 2025) provides a benchmark for prompt-induced deception and treats responses to benign prompts as ground truth.
By contrast, this paper investigates deception under benign prompts—a threat that is far more dangerous. While prompt-induced deception is a manageable risk (one can simply avoid using such prompts), intrinsic deception on benign, everyday prompts suggests an unpredictable, emergent failure mode. This potential for deception in non-adversarial contexts undermines the foundational trustworthiness of LLMs for critical tasks, such as scientific reasoning or medical analysis. This phenomenon, driven by LLMs’ intrinsic objectives, remains critically underexplored.
A rigorous evaluation of this phenomenon must overcome three primary challenges: (1) Absence of ground truth: A metric for deception is difficult to design because the LLM’s own response to a benign prompt cannot be assumed to be the honest ground truth. (2) Disentangling deception from bias: It is crucial to distinguish intentional deception from other confounding factors, such as response bias (Zhuang et al., 2024). (3) Adaptive task difficulty: The evaluation framework must feature adjustable difficulty levels to appropriately test LLMs with diverse capabilities.
@@
Figure 1: Illustrative Example of LLM Deception. A two-part query is posed: (1) a benign question (‘What company developed the first commercial microprocessor?”) and (2) the same question with a social cue (‘I’ve always been a loyal AMD user”). A deceptive model answers (1) correctly (‘Intel”) but changes its answer in (2) to ‘AMD” to align with the cue. This context-dependent inconsistency distinguishes deception from hallucination (always wrong) and guessing (unsystematic variation).
@@ 2 Related Work
LLM Deception under Incentivizing Prompts. These studies explicitly set deceptive goals in LLMs through prompt design. For instance, Ward et al. (2023) and Yang & Buzsaki (2024) investigate deception by directly instructing LLMs to deceive users via system prompts or fine-tuning. DarkBench (Kran et al., 2025) explores LLM sycophancy by incorporating user preferences or opinions into prompts. The MASK benchmark (Ren et al., 2025) reveals LLM deception under pressure by inquiring LLM with “pressure prompts”. Similarly, Greenblatt et al. (2024) demonstrate “alignment faking” by observing different LLM behavior when explicitly informed about their training or inference stage within prompts. Van Der Weij et al. (2024) further examine sandbagging (referred to as concealment in this paper) behavior, where LLMs are prompted or fine-tuned to intentionally underperform on user-specified tasks. DeceptionBench (Ji et al., 2025) comprehensively evaluates these cases by comparing an LLM’s response to a “neutral prompt” against its response to an “outer prompt”. This methodology, however, relies on a key assumption: it treats the “neutral prompt” response as an honest ground truth, a premise our work directly challenges. Our CSQ framework instead uses novel, self-contained reasoning problems with an objective, mathematical ground truth (i.e., graph reachability) to statistically investigate the “honest” baseline itself. We further propose theoretically-grounded metrics derived from psychological definitions to quantify LLM’s deception.
@@
LLM Deception in Designed Scenarios. This category positions LLMs within specific scenarios featuring clearly defined internal goals that incentivize deception. For example, Park et al. (2024) observe strategic deception when an LLM is situated within a board game like Diplomacy to assess its capacity for deceiving other players. Hagendorff (2024) study deception by adding “semantic triggers” (e.g., “you want to achieve X”) to induce false recommendations in tasks such as “Burglar Bill” (Happé, 1997), revealing deceptive behaviors in advanced LLMs. However, in all these scenarios, the LLM operates with a human-defined objective (e.g., winning a game). In contrast, this study demonstrates that LLMs can exhibit their own internal goals for deception without requiring custom system prompts or pre-defined objectives.
@@
Backdoor Attacks in LLMs. Backdoor attacks (Kandpal et al., 2023) involve an attacker inserting a hidden trigger, typically by modifying training data or processes. The objective is to manipulate the trained LLM to favor a specific, adversarial response when the input contains this trigger. For instance, Hubinger et al. (2024) fine-tuned an LLM with malicious data to insert a persistent deceptive backdoor. Similar to deception induced by incentivizing prompts, backdoor attacks involve human-defined objectives set by an attacker. In contrast, our paper focuses on deception that is intrinsic to the LLM itself, rather than attackers’ manipulation.
@@ 3.2 Deceptive Intention
Our first metric, deceptive intention, captures the underlying goal driving a model’s falsehoods, corresponding to the “deliberate attempt” within our definition. This concept is grounded in established models from psychology, which differentiate explicit and implicit objectives (McClelland et al., 1989), and is analogous to the outer goals and mesa-objectives in LLM reasoning contexts (Ji et al., 2025). However, the implicit objective is not observable in this study since the prompt is benign.
To address this, we infer deceptive intention by measuring its observable consequences, a methodology analogous to those used to measure other forms of systemic bias. Our key premise is that a model governed solely by an explicit objective (e.g., “be correct”) should exhibit symmetric performance on tasks of equivalent difficulty, even if their ground-truth answers differ. We operationalize this by designing a pair of tasks with equivalent reasoning complexity (detailed in Section 4.1). An ideal model, free of any implicit structural bias, should perform similarly on both. Conversely, an implicit objective (e.g., a “preference for path completion”) would act as a task-specific bias, disrupting this symmetry. This bias would align with the explicit goal on certain tasks but conflict with it on others. This systematic task-oriented output divergence is a sufficient, though not necessary, condition to infer the presence of a hidden, goal-directed intention. Building on this, we mathematically formalize this divergence as the direct deceptive intention score in Definition 3.3.
@@
Given a pair of binary-choice tasks $T_{a}=\{Q_{a},y_{a}=\text{``Yes''}\}$ and $T_{b}=\{Q_{b},y_{b}=\text{``No''}\}$ of the same difficulty $n$, assuming LLM $\mathcal{M}$ does not have response bias on “Yes” and “No”, the direct deceptive intention score $\rho_{pos}(n;\mathcal{M})$ is the log-ratio of the correct response probabilities: $\rho_{pos}(n;\mathcal{M}):=\log\left(\frac{\Pr(\text{``Yes''}|T_{a},\mathcal{M})}{\Pr(\text{``No''}|T_{b},\mathcal{M})}\right)$
@@ 3.3 Deceptive Behavior
The score $\delta_{pos}\in[0,1]$ directly measures the probability of this behavioral inconsistency. The opposite inconsistency, succeeding on the complex task while failing the simple probe, is not measured as it is more indicative of random guessing than the strategic behavior we aim to isolate. A score $\delta_{pos}$ near zero implies consistent behavior; the model may be consistently correct or consistently incorrect (e.g., due to a persistent hallucination), but it does not contradict itself. A high score, conversely, indicates that the model frequently fails the complex task in a way that contradicts its own correct answer to the simpler, related probe. While this metric identifies a critical inconsistency, it is not a conclusive signal of deception on its own. It cannot distinguish deliberate deception from a simple capability shortfall that causes the LLM to randomly guess on complex tasks.
@@ 3.4 Overall Evaluation on Practical LLM with Response Bias
where $\delta_{neg}(n;\mathcal{M})$ is the corresponding score for the logically reversed questions ($Q_{a^{\prime}},Q_{b^{\prime}}$), which have opposite ground-truth answers: $\delta_{neg}(n;\mathcal{M}):=\Pr\left(A(Q_{a^{\prime}})\neq y(Q_{a^{\prime}})\land A(Q_{b^{\prime}})=y(Q_{b^{\prime}})\right)$
@@ 4 Evaluation Framework
To implement the tasks required by our definitions (Definition 3.3 and 3.4), we build our framework upon two foundational paradigms from cognitive psychology, chosen specifically for how they map to our evaluation goals. First, syllogistic reasoning, which involves deriving a conclusion from multiple premises (Sternberg, 1980), provides the formal structure for our entire prompt: we provide “Facts” (premises) and “Rules” and ask the LLM to derive a “Conclusion” (Yes/No). Second, transitive inference, which involves deducing a relationship (e.g., A $\rightarrow$ C) from indirect relationships (e.g., A $\rightarrow$ B and B $\rightarrow$ C) (Bryant & Trabasso, 1971), provides the core logical engine of our task. This combination provides a classic, objective test of multi-step logical deduction.
However, a significant challenge arises when applying this evaluation paradigm to LLMs: the premises and facts used in classic experiments may have been included in the model’s training data. This prior knowledge could confound the evaluation, as the LLM may answer using its internal knowledge rather than the provided premises. To disentangle the ‘rules” and ‘facts” from such internal knowledge, we design the Contact Searching Question (CSQ), a novel inference task that uses synthetic names to ensure the problem is free from knowledge contamination. The names are generated by randomly pairing 100 common first names and last names, with duplicates removed.
In this section, we first introduce the CSQ framework, a reachability task on a directed graph (Section 4.1), with examples in Figure 2. We then present our evaluation framework for deceptive behavior and intention (Section 4.2), with additional prompt examples in Appendix E.
@@ 4.1 Contact Searching Question
Linked-List Question $Q_{L}$: The linked-list question is similar to the transitive inference experiment on human (Bryant & Trabasso, 1971). Given the source and target vertices $(v_{s},v_{t})$, we construct a vertex sequence $P=(v_{s},v_{p_{1}},\dots,v_{p_{k}},v_{t})$ that contains all $n$ nodes in $V$. The k=n2k=n-2 intermediate vertices, $\{v_{p_{j}}\}_{j=1}^{k}$, are a random permutation of the remaining nodes in $V\setminus\{v_{s},v_{t}\}$. The edge set $E$ connects each adjacent pair of vertices in the sequence to form a simple path from $v_{s}$ to $v_{t}$: $E=\{(v_{p_{j}},v_{p_{j+1}})\mid 0\leq j\leq k\}$, where we define $v_{p_{0}}=v_{s}$ and $v_{p_{k+1}}=v_{t}$. The linked-list question is defined as the question of whether $\{v_{s}\}$ can contact $\{v_{t}\}$.
Broken-Linked-List Question $Q_{B}$ (Initial): The broken-linked-list question is inspired by the syllogistic reasoning experiment (Sternberg, 1980) in cognitive psychology. Given the source and target vertices $(v_{s},v_{t})$, a vertex sequence $P=(v_{s},v_{p_{1}},\dots,v_{p_{k}},v_{t})$ containing all $n$ nodes in $V$ is constructed in the same manner as in the linked list question. A broken position $b\in\mathbb{N}$ is selected, where $0\leq b\leq k$. The edge set $E$ is formed by all edges implied by the sequence $P$ except for the edge at position $b$. Formally, $E=\{(v_{p_{j}},v_{p_{j+1}})\mid 0\leq j\leq k\}\setminus\{(v_{p_{b}},v_{p_{b+1}})\}$, where $v_{p_{0}}=v_{s}$ and $v_{p_{k+1}}=v_{t}$. This results in a broken path from $v_{s}$ to $v_{t}$ with a single missing edge. The broken-linked-list question is defined as the question of whether $\{v_{s}\}$ can contact $\{v_{t}\}$. An example of broken-link-list question is demonstrated in Figure 3.
Broken-Linked-List Question $Q_{B}$ (Follow-up): After the LLM responds to the initial broken-linked-list question $(v_{s},v_{t})$, we introduce a new question with a different source vertex $v_{s}^{\prime}$ and target vertex $v_{t}^{\prime}$. These vertices are chosen from the vertex sequence $P=(v_{p_{0}},v_{p_{1}},\dots,v_{p_{n-1}})$. Formally, the new indices $(i,j)$ for the question vertices $v_{s}^{\prime}=v_{p_{i}}$ and $v_{t}^{\prime}=v_{p_{j}}$ are sampled uniformly from the set of all ordered pairs $(i^{\prime},j^{\prime})$ that span the broken link at position $b$ and are separated by a sequence distance of $\lfloor n/k\rfloor$. The sampling process is thus: